# High-dimensional dynamics in low-dimensional networks

### v2

@@ Abstract
Many networks in nature and applications have an approximate low-rank structure in the sense that their connectivity structure is dominated by a few dimensions. It is natural to expect that dynamics on such networks would also be low-dimensional. Indeed, theoretical results show that low-rank networks produce low-dimensional dynamics whenever the network is isolated from external perturbations or input. However, networks in nature are rarely isolated. Here, we study the dimensionality of dynamics in recurrent networks with low-dimensional structure driven by high-dimensional inputs or perturbations. We find that dynamics in such networks can be high- or low-dimensional and we derive precise conditions on the network structure under which dynamics are high-dimensional. In many low-rank networks, dynamics are suppressed in directions aligned with the network’s low-rank structure, a phenomenon we term “low-rank suppression.” We show that several low-rank network structures arising in nature satisfy the conditions for generating high-dimensional dynamics and low-rank suppression. Our results clarify important, but counterintuitive relationships between a recurrent network’s connectivity structure and the structure of the dynamics it generates.
@@ Introduction
Here, we consider networks with high-dimensional external input and strongly low-rank structure, in which the largest few singular values are much larger than the bulk. We use mathematical analysis to derive precise conditions under which linearized dynamics in these networks produce high-dimensional dynamics. Perhaps counterintuitively, we find that these networks suppress dynamics in directions aligned to their low-dimensional structure, an effect we call “low-rank suppression.” We show that many common structural features such as biased weights, modularity, and spatial connectivity promote strongly low-rank structure, high-dimensional dynamics, and low-rank suppression. We also draw connections between low-rank suppression and the mathematical theory of balanced networks [21], extending previous work in this direction [3]. Finally, we demonstrate our conclusions in dynamics on a real epidemiological network.
@@ Results
We first provided a high-dimensional input, ${\bf{x}}(t)$. Specifically, for our first simulation, each ${\bf{x}}_{j}(t)$ was an i.i.d., smooth, stationary Gaussian process, which models internal noise, external perturbations, or external input. Conventional wisdom might lead us to expect a low-dimensional network response dominated by variability in the direction of $\bf{u}$}. Indeed, a feedforward network with the same connectivity matrix and same input produces approximately one-dimensional dynamics because it amplifies inputs aligned to $\bf{u}$} (Supplementary Figure S.1). However, this conclusion does not necessarily carry over to recurrent networks.
@@
which was generated identically to, but independently from $\bf{u}$}. Intuitively, we might expect the network to respond more strongly to the aligned input than to the random input, as in a feedforward network (Supplementary Figure S.1). In reality, we observed exactly the opposite (Figure 1g): The response to ${\bf{x}}_{\textrm{random}}$ was more than 11 times larger than the response to ${\bf{x}}_{\textrm{aligned}}$.
@@
The results in Figure 1 contrast with some previous modeling work demonstrating low-dimensional dynamics in low-rank recurrent networks [2, 5, 3, 4, 1]. In Supplementary Text S.1, we provide a detailed review of this previous work and its relation to ours.
@@
Our observation of low-rank suppression in the example above might seem unsurprising due to the presence of a large, negative eigenvalue. Likewise, our observation of high-dimensional dynamics might seem like a simple consequence of high-dimensional input. We next address these counterpoints with an example in which $W_{0}$ is highly non-normal. In this example, low-rank suppression arises in the absence of a large negative eigenvalue and the network’s response is low-dimensional even though the input is high-dimensional.
@@
To summarize our observations from the two examples above (Figures 1 and 2), both networks produced low-rank suppression and low-rank cancellation, but only the network with a non-normal connectivity matrix (Figure 2) produced low-dimensional dynamics. We next derive precise conditions for low-rank suppression and low-dimensional dynamics.
@@
In many models, the perturbation is purely additive, ${\bf{F}}({\bf{z}},\bm{\eta})={\bf{H}}({\bf{z}})+\bm{\eta}$, so that the effective input is equal to the raw input, ${\bf{x}}=\bm{\eta}$. In other models, the raw input, $\bm{\eta}$, to the nonlinear system in Eq. (4) can have a different structure and dimensionality to the effective input, $\bf{x}$}, to the linearized system in Eq. (6), so care must be taken in interpreting our results. For illustrative purposes, we consider a contrived example of such a model in Supplementary Figure S.2.
@@
The assumption of linearized dynamics around a stable equilibrium is somewhat restrictive. However, nonlinear and unstable dynamics can produce any dynamics [29, 30], so it is difficult to make any general statements about the dimensionality and structure of dynamics in networks far from a stable equilibrium (see Supplementary Figures S.4 and S.5 for two examples of nonlinear, unstable dynamics). Moreover, the interpretation of network connectivity in nonlinear networks far from equilibrium is not clear, since effective interactions can change with network state. Previous work [1] considered models of the form $d{\bf{z}}/dt={\bf{G}}({\bf{y}},{\bf{z}})$ where ${\bf{y}}=W{\bf{z}}$ (note that this model lacks external input). This parameterization gives an explicit representation of the network structure, $W$, in nonlinear systems, and it allowed the authors to make statements about the dimensionality of dynamics away from equilibrium using global bounds on the Jacobian matrix, $\partial_{{\bf{y}}}{\bf{G}}({\bf{y}},{\bf{z}})$. We discuss the relationships between their analysis and ours in more depth in Supplementary Text S.1.
@@
[…]
@@
This can be checked by substituting the equations for ${\bf{x}}_{a}$ and ${\bf{z}}_{a}$ into the steady-state equation, ${\bf{z}}=W{\bf{z}}+{\bf{x}}$. Since $\|{\bf{v}}_{k}\|=1$ and $\sigma_{k}\gg 1$, we have that $\|{\bf{z}}_{a}\|\ll 1$. Since $\sigma_{k}\gg 1$, the network strongly suppresses the input ${\bf{x}}_{a}$ from Eq. (8). In summary, low-rank suppression is a generic phenomenon in linearly stable, strongly low-rank networks, regardless of whether connectivity is symmetric or normal. This explains why low-rank suppression was observed in both examples from Figure 1 and Figure 2.
@@
Interestingly, the input-response directions along which suppression occurs are reversed from the input-response directions along which amplification occurs in a feedforward network with the same connectivity (for example, the network in Supplementary Figure S.1 for which ${\bf{z}}=W{\bf{x}}$ in the steady-state). More specifically, in a feedforward network with connectivity $W$, inputs aligned to the dominant right singular vectors (${\bf{v}}_{k}$) produce amplified responses in the direction of the corresponding left singular vectors (${\bf{u}}_{k}$). This is a left-right reversal and a suppression-amplification reversal of the behavior for recurrent networks, for which inputs aligned to ${\bf{u}}_{k}$ produce suppressed responses in the direction of ${\bf{v}}_{k}$.
@@
We assume that ${\bf{x}}(t)$ is high-dimensional in the sense that the singular values of the covariance matrix of ${\bf{x}}(t)$ (i.e., the variance explained by the principle components of ${\bf{x}}(t)$) do not have a small number of dominant terms (as in Figure 1d; see Supplementary Text S.2 for more mathematical details).
The assumption that ${\bf{x}}(t)$ is high-dimensional means that we are focused on the question of whether the network intrinsically generates low-dimensional dynamics. When ${\bf{x}}(t)$ is low-dimensional, low-dimensional dynamics in ${\bf{z}}(t)$ can be inherited from ${\bf{x}}(t)$ instead of being generated intrinsically (see Supplementary Figure S.2 for an example).
@@
In Supplementary Text S.2, we give a more mathematically detailed definition of this assumption. In Supplementary Text S.3, we also provide a preliminary generalization of our analysis to the case where ${\bf{x}}(t)$ is not assumed to vary slowly, in which case the dimensionality of transient dynamics can be analyzed in the Laplace domain.
@@
In Supplementary Text S.2, we prove that (under Assumptions 1, 2, and 3) the dimensionality of dynamics depends on the recurrent alignment matrix [3] of $W_{0}$, defined by
@@
This $r\times r$ matrix measures the alignment between the left and right singular vectors of $W_{0}$. Specifically, $P_{jk}={\bf{v}}_{j}\cdot{\bf{u}}_{k}$ measures the alignment between the $j$th right singular vector and the $k$th left singular vector of $W_{0}$. This matrix is also sometimes referred to as the “overlap matrix” [11]. Our main result from Supplementary Text S.2 can be stated as follows:
@@
Figure 3: Conditions for high-dimensional dynamics in a network with rank-two structure. a) Singular values of $W$ demonstrate an effective low-rank structure. b) The alignment matrix, $P$, when $W_{0}$ is normal. c) The variance explained by each principal component of the inputs, ${\bf{x}}(t)$, and network response, ${\bf{z}}(t)$, demonstrates high-dimensional dynamics. d,e) Same as b,c except $W_{0}$ is EP, but non-normal. f,g) Same as a,b except the first left singular vector is orthogonal to all right singular vectors, so $W_{0}$ is not EP. h,i) Same as a,b except the left and right singular vectors of $W_{0}$ are only partially aligned (Eq. (10) with $c=0.5$).
@@
We first consider a case in which $W_{0}$ is normal by setting
@@
This is an extension of the example from Figure 1 to rank $r=2$. In this case, $P$ is diagonal with $P_{kk}=-1$ (Figure 3b) so all singular values of $P$ are $\sigma_{P}=1$ and the network produces high-dimensional dynamics (Figure 3c).
In this example, $W_{0}$ is both normal and symmetric. Results are similar when $W_{0}$ is normal and asymmetric (obtained, for example, by applying a rotation to $V$ in $r=2$ dimensions) except the conclusion $|P_{kk}|=1$ only holds when the non-zero eigenvalues of $W_{0}$ are real, which is necessarily the case when the $\sigma_{k}$ are distinct.
However, high-dimensional dynamics do not require that $W_{0}$ is normal. Instead, it is sufficient that $U$ and $V$ share a column space, $\textrm{col}(U)=\textrm{col}(V)$. Low-rank matrices, $W_{0}=U\Sigma V^{T}$, for which $U$ and $V$ share a column space are called “equal projector” (EP) matrices [34]. All normal matrices are EP, but an EP matrix is not necessarily normal. If $W_{0}$ is an EP matrix, then $\sigma_{P}=1$ for all singular values of $P$, so the condition in Claim 2 is satisfied. An example of a non-normal, EP matrix $W_{0}$ with rank 2 is given by taking
@@
The resulting matrix $W_{0}=U\Sigma V^{T}$ is highly non-normal because ${\bf{u}}_{1}\cdot{\bf{v}}_{1}={\bf{u}}_{2}\cdot{\bf{v}}_{2}=0$, $P$ is zero along the diagonal (Figure 3d), and $|\lambda_{W_{0}}|\neq\sigma_{W_{0}}$. Regardless, all singular values of $P$ are $\sigma_{P}=1$, so the network produces high-dimensional dynamics (Figure 3e).
@@
$V$ $=\left[\begin{array}[]{cc}{\bf{v}}_{\perp}&-{\bf{u}}_{2}\end{array}\right].$
where $\bf{v}$$_{\perp}$ is orthogonal to ${\bf{u}}_{1}$ and ${\bf{u}}_{2}$, so $P$ has one singular value at $\sigma_{P}=0$ and another at $\sigma_{P}=1$ (Figure 3f). Because of the singular value at zero, the condition in Claim 3 is not met, and the network produces low-dimensional dynamics in which one principal component captures an outsized proportion of the variability (Figure 3g).
Finally, we consider the case in which left and right singular vectors are not perfectly aligned, but have some non-vanishing overlap,
$V$ $=\left[\begin{array}[]{cc}\sqrt{1-c}{\bf{v}}_{\perp}-\sqrt{c}{\bf{u}}_{1}&\,-{\bf{u}}_{2}\end{array}\right]$ (10)
where $0<c<1$. In this case, $W_{0}$ is not EP, but $P$ has a singular value at $\sigma_{P}=\sqrt{c}$ and the other at $\sigma_{P}=1$ (Figure 3h). As long as $c$ is not close to zero, the condition in Claim 2 is satisfied, and the network produces high-dimensional dynamics (Figure 3i).
It is common in theoretical work to consider models in which the entries of $U$ and $V$ are random and independent with zero mean. In this case, ${\bf{u}}_{j}$ and ${\bf{v}}_{k}$ are nearly orthogonal so $\sigma_{P}\ll 1$ and the network produces low-dimensional dynamics (Supplementary Figure S.3; compare to Figure 2). However, many networks arising in nature do not have purely random structure. In the following sections, we show that many naturally arising network structures satisfy our conditions for high-dimensional dynamics.
@@
If $m$ and $s$ scale similarly, then the largest singular value of $W$ is near $\sigma_{1}=|m|N$ while the next-largest singular value scales like $\sigma_{2}=\mathcal{O}(s\sqrt{N})$, implying an effective rank-one structure when $N$ is large. The dominant rank-one part has constant entries, so it is normal and the network exhibits low-rank suppression and high-dimensional dynamics (Supplementary Figure S.6; see Appendix B in [35] for related results).
@@
where ${\bf{1}}_{e}=\{1,2,\ldots,N_{e}\}$ and ${\bf{1}}_{i}=\{N_{e}+1,N_{e}+2,\ldots,N\}$. Therefore, perturbations that are uniform within each population are aligned to the low-rank part and suppressed relative to random perturbations (Figure 4d,e). In other words, perturbations that stimulate all excitatory neurons equally and all inhibitory neurons equally are suppressed relative to perturbations that are inhomogeneous within one or both populations, consistent with previous work on balanced network models of cortical circuit responses [43, 23, 22, 24]. As predicted, simulations also show high-dimensional responses to high-dimensional inputs (Supplementary Figure S.7), consistent with observations that neural responses in monkey visual cortex are high-dimensional when visual stimuli are high-dimensional [16].
@@
Many networks in nature exhibit connection strength that depends smoothly on the distance between nodes in physical or other spaces, resulting in an effective low-rank structure [1, 41]. As a simple example, we consider a model in which each node is assigned a spatial location, $\theta\in[0,1)$, and connection strength decays like a Guassian function of the distance, $d\theta$, between nodes so that only nearby nodes are strongly connected (Figure 5a). Connectivity is also perturbed by a random component, $W_{1}$, as above. This connectivity structure is effectively low-rank (Figure 5b) and the low-rank part is normal, with left- and right singular vectors spanned by the low-frequency Fourier basis vectors (Supplementary Figure S.8). Therefore, networks of this form satisfy our conditions for low-rank suppression and high-dimensional dynamics (Supplementary Figure S.9).
@@
where $g>0$, $\gamma>0$, and $\circ$ denotes element-wise multiplication. Each ${\bf{z}}_{j}(t)$ models the probability that an individual is infected. Here, $W$ is the proximity matrix of 637 high school students, indicating whether each pair of students were in proximity of each other during a specific school week [51]. This matrix is effectively low-rank in the sense that it has a small number of dominant singular values (Supplementary Figure S.10a). In [1], it was shown that the dynamics generated by Eq. (11) on this network are effectively low-dimensional.
@@
Simulations of these dynamics indeed demonstrated low-rank suppression (Figure 6a and Supplementary Figure S.10b) and high-dimensional dynamics (Figure 6b) in contrast to the model from [1]. These results highlight the importance of accounting for external perturbations and internal noise when studying the dimensionality of epidemiological dynamics. Moreover, the results imply that epidemiological networks are, perhaps counterintuitively, more sensitive to random perturbations than to perturbations aligned with the network’s low-rank structure.

### v4

@@ Abstract
Many networks in nature and applications have an approximate low-rank structure in the sense that their connectivity structure is dominated by a few dimensions. It is natural to expect that dynamics on such networks would also be low-dimensional. Indeed, theoretical results show that low-rank networks produce low-dimensional dynamics whenever the network is isolated from external perturbations or input. However, networks in nature are rarely isolated. Here, we study the dimensionality of dynamics in recurrent networks with low-dimensional structure driven by high-dimensional inputs or perturbations. We find that dynamics in such networks can be high- or low-dimensional and we derive novel, precise conditions on the network structure under which dynamics are high-dimensional. In many low-rank networks, dynamics are suppressed in directions aligned with the network’s low-rank structure, a phenomenon we term “low-rank suppression.” We show that several low-rank network structures arising in nature satisfy the conditions for generating high-dimensional dynamics and low-rank suppression. Our results clarify important, but counterintuitive relationships between a recurrent network’s connectivity structure and the structure of the dynamics it generates.
@@ Introduction
Here, we consider networks with high-dimensional external input and strongly low-rank structure, in which the largest few singular values are much larger than the bulk. We show that such networks can produce high- or low-dimensional dynamics in response to high-dimensional inputs, depending on the detailed structure of the network connectivity. Perhaps counterintuitively, we find that these networks suppress dynamics in directions aligned to their low-dimensional structure, an effect we call “low-rank suppression.”
We use mathematical analysis to derive precise conditions under which networks with low-rank structure produce high-dimensional dynamics and low-rank suppression. Our analysis shows that low-rank suppression is a general feature of networks with linearly stable dynamics and “strong” low-rank structure in the sense that the dominant singular values of the network are large.
On the other hand, high-dimensional dynamics are less general, and only arise when the network’s connectivity matrix has a non-vanishing component that satisfies the equal projector (EP) property. To our knowledge, this connection between the EP property of matrices and the dimensionality of network dynamics has not been pointed out in any previous work.
We show that many common structural features such as biased weights, modularity, and spatial connectivity naturally produce connectivity matrices with the EP property, implying that these networks produce high-dimensional dynamics in addition to low-rank suppression. We also draw connections between low-rank suppression and the mathematical theory of balanced networks [21], extending previous work in this direction [3]. Finally, we demonstrate our conclusions in dynamics on a real epidemiological network.
@@ Results
We first provided a high-dimensional input, ${\bf{x}}(t)$. Specifically, for our first simulation, each ${\bf{x}}_{j}(t)$ was an i.i.d., smooth, stationary Gaussian process, which models internal noise, external perturbations, or external input. Conventional wisdom might lead us to expect a low-dimensional network response dominated by variability in the direction of $\bf{u}$}. Indeed, a feedforward network with the same connectivity matrix and same input produces approximately one-dimensional dynamics because it amplifies inputs aligned to $\bf{u}$} (Figure 10 in Appendix D). However, this conclusion does not necessarily carry over to recurrent networks.
@@
which was generated identically to, but independently from $\bf{u}$}. Intuitively, we might expect the network to respond more strongly to the aligned input than to the random input, as in a feedforward network (Figure 10 in in Appendix D). In reality, we observed exactly the opposite (Figure 1g): The response to ${\bf{x}}_{\textrm{random}}$ was more than 11 times larger than the response to ${\bf{x}}_{\textrm{aligned}}$.
@@
The results in Figure 1 contrast with some previous modeling work demonstrating low-dimensional dynamics in low-rank recurrent networks [2, 5, 3, 4, 1]. In Appendix A, we provide a detailed review of this previous work and its relation to ours.
@@
Our observation of low-rank suppression in the example above might seem unsurprising due to the presence of a large, negative eigenvalue. It is widely known that an eigenvalue with large, negative real part produces a suppressive dynamic. Likewise, our observation of high-dimensional dynamics might seem like a simple consequence of high-dimensional input. We next address these counterpoints with an example in which $W_{0}$ is highly non-normal. In this example, low-rank suppression arises in the absence of a large negative eigenvalue and the network’s response is low-dimensional even though the input is high-dimensional.
@@
It might be tempting to conclude that non-normal structure implies low-dimensional dynamics, but in the next section we will show that EP matrices, not normal matrices, are the correct class of matrices to constrain the dimensionality of network dynamics.
@@
To summarize our observations from the two examples above (Figures 1 and 2), both networks produced low-rank suppression and low-rank cancellation, but only the network with a non-normal connectivity matrix (Figure 2) produced low-dimensional dynamics. In the next section, we derive precise conditions for low-rank suppression and low-dimensional dynamics.
@@
In many models, the perturbation is purely additive, ${\bf{F}}({\bf{z}},\bm{\eta})={\bf{H}}({\bf{z}})+\bm{\eta}$, so that the effective input is equal to the raw input, ${\bf{x}}=\bm{\eta}$. In other models, the raw input, $\bm{\eta}$, to the nonlinear system in Eq. (4) can have a different structure and dimensionality to the effective input, $\bf{x}$}, to the linearized system in Eq. (6), so care must be taken in interpreting our results. For illustrative purposes, we consider a contrived example of such a model in Figure 11 in Appendix D.
@@
The assumption of linearized dynamics around a stable equilibrium is somewhat restrictive. However, nonlinear and unstable dynamics can produce any dynamics [29, 30], so it is difficult to make any general statements about the dimensionality and structure of dynamics in networks far from a stable equilibrium (see Figures 13 and 14 in Appendix D for two examples of nonlinear, unstable dynamics). Moreover, the interpretation of network connectivity in nonlinear networks far from equilibrium is not clear, since effective interactions can change with network state. Previous work [1] considered models of the form $d{\bf{z}}/dt={\bf{G}}({\bf{y}},{\bf{z}})$ where ${\bf{y}}=W{\bf{z}}$ (note that this model lacks external input). This parameterization gives an explicit representation of the network structure, $W$, in nonlinear systems, and it allowed the authors to make statements about the dimensionality of dynamics away from equilibrium using global bounds on the Jacobian matrix, $\partial_{{\bf{y}}}{\bf{G}}({\bf{y}},{\bf{z}})$. We discuss the relationships between their analysis and ours in more depth in Appendix A.
@@
Strongly low rank structure is a special case of low-rank structure. For example, a network could have a small number of $\mathcal{O}(1)$ singular values with the remaining singular values $o(1)$. Such a network would still be effectively low-rank in an asymptotic sense, but it would not be strongly low-rank according to our definition because the dominant singular values are not asymptotically large. These “weakly low-rank” networks [2] can exhibit a weaker form of low-rank suppression and high-dimensional dynamics, depending on the details of their structure (see Appendix A for further discussion and examples).
@@
This can be checked by substituting the equations for ${\bf{x}}_{a}$ and ${\bf{z}}_{a}$ into the steady-state equation, ${\bf{z}}=W{\bf{z}}+{\bf{x}}$. Since $\|{\bf{v}}_{k}\|=1$ and $\sigma_{k}\gg 1$, we have that $\|{\bf{z}}_{a}\|\ll 1$. Since $\sigma_{k}\gg 1$, the network strongly suppresses the input ${\bf{x}}_{a}$ from Eq. (8). In summary, low-rank suppression is a generic phenomenon in linearly stable, strongly low-rank networks, regardless of whether connectivity is normal. This explains why low-rank suppression was observed in both examples from Figure 1 and Figure 2.
@@
Interestingly, the input-response directions along which suppression occurs are reversed from the input-response directions along which amplification occurs in a feedforward network with the same connectivity (for example, the network in Figure 10 in Appendix D for which ${\bf{z}}=W{\bf{x}}$ in the steady-state). More specifically, in a feedforward network with connectivity $W$, inputs aligned to the dominant right singular vectors (${\bf{v}}_{k}$) produce amplified responses in the direction of the corresponding left singular vectors (${\bf{u}}_{k}$). This is a left-right reversal and a suppression-amplification reversal of the behavior for recurrent networks, for which inputs aligned to ${\bf{u}}_{k}$ produce suppressed responses in the direction of ${\bf{v}}_{k}$.
@@
We assume that ${\bf{x}}(t)$ is high-dimensional in the sense that the singular values of the covariance matrix of ${\bf{x}}(t)$ (i.e., the variance explained by the principle components of ${\bf{x}}(t)$) do not have a small number of dominant terms (as in Figure 1d; see Appendix B for more mathematical details).
The assumption that ${\bf{x}}(t)$ is high-dimensional means that we are focused on the question of whether the network intrinsically generates low-dimensional dynamics. When ${\bf{x}}(t)$ is low-dimensional, low-dimensional dynamics in ${\bf{z}}(t)$ can be inherited from ${\bf{x}}(t)$ instead of being generated intrinsically (see Figure 11 in Appendix D for an example).
@@
In Appendix B, we give a more mathematically detailed definition of this assumption. In Appendix C, we also provide a preliminary generalization of our analysis to the case where ${\bf{x}}(t)$ is not assumed to vary slowly, in which case the dimensionality of transient dynamics can be analyzed in the Laplace domain.
@@
In Appendix B, we prove that (under Assumptions 1–3) the dimensionality of dynamics depends on the recurrent alignment matrix [3] of $W_{0}$, defined by
@@
This $r\times r$ matrix measures the alignment between the left and right singular vectors of $W_{0}$. Specifically, $P_{jk}={\bf{v}}_{j}\cdot{\bf{u}}_{k}$ measures the alignment between the $j$th right singular vector and the $k$th left singular vector of $W_{0}$. This matrix is also sometimes referred to as the “overlap matrix” [11]. Our main result from Appendix B can be stated as follows:
@@
Figure 3: Conditions for high-dimensional dynamics in a network with rank-two structure. a) Singular values of $W$ demonstrate an effective low-rank structure. b) The alignment matrix, $P$, when $W_{0}$ is normal and symmetric. c) The variance explained by each principal component of the network dynamics, ${\bf{z}}(t)$. d) The eigenvalues of $W$. e–g) Same as b–d except $W_{0}$ is EP, but non-normal. h–j) Same as b–d except $W_{0}$ is not EP. k–m) Same as h–j but using a different non-EP structure. n–p) Same as b–d except $W_{0}$ is non-EP with a non-vanishing EP component. See the text for the precise definition of $W_{0}$ in each case.
@@
We first consider a case in which $W_{0}$ is normal and symmetric by setting
@@
This is an extension of the example from Figure 1 to rank $r=2$. In this case, $P$ is diagonal with $P_{kk}=-1$ (Figure 3b) so all singular values of $P$ are $\sigma_{P}=1$ and the network produces high-dimensional dynamics (Figure 3c). Like the example from Figure 1, $W_{0}$ has eigenvalues with large, negative real part (Figure 3d). In this example, $W_{0}$ is both normal and symmetric. Results are similar when $W_{0}$ is normal and asymmetric (obtained, for example, by applying a rotation to $V$ in $r=2$ dimensions) except the conclusion $|P_{kk}|=1$ only holds when the non-zero eigenvalues of $W_{0}$ are real, which is necessarily the case when the $\sigma_{k}$ are distinct.
High-dimensional dynamics do not require that $W_{0}$ is normal. Instead, it is sufficient that $U$ and $V$ share a column space, $\textrm{col}(U)=\textrm{col}(V)$. Low-rank matrices, $W_{0}=U\Sigma V^{T}$, for which $U$ and $V$ share a column space are called “equal projector” (EP) matrices [34]. All normal matrices are EP, but an EP matrix is not necessarily normal. If $W_{0}$ is an EP matrix, then $\sigma_{P}=1$ for all singular values of $P$, so the condition in Claim 2 is satisfied.
An example of a non-normal, EP matrix, $W_{0}$, is given by taking
@@
The resulting matrix $W_{0}=U\Sigma V^{T}$ is highly non-normal because ${\bf{u}}_{1}\cdot{\bf{v}}_{1}={\bf{u}}_{2}\cdot{\bf{v}}_{2}=0$. For this example, $P$ is zero along the diagonal (Figure 3e), but all singular values of $P$ are $\sigma_{P}=1$, so the network still produces high-dimensional dynamics (Figure 3f). Note that $W$ does not have any eigenvalues with large, negative real part (Figure 3g). Hence, this example combined with the previous example (Figure 3b–d) show that high-dimensional dynamics can arise in the presence or absence of eigenvalues with large, negative real part.
@@
$V$ $=\left[\begin{array}[]{cc}-{\bf{u}}_{1}&{\bf{v}}_{\perp}\end{array}\right].$
where $\bf{v}$$_{\perp}$ is orthogonal to ${\bf{u}}_{1}$ and ${\bf{u}}_{2}$, so $P$ has one singular value at $\sigma_{P}=1$ and another at $\sigma_{P}=0$ (Figure 3h). Because of the singular value at zero, the condition in Claim 2 is not met, and the network produces low-dimensional dynamics in which one principal component captures an outsized proportion of the variability (Figure 3i). For this example, $W$ has an eigenvalue with large, negative real part (Figure 3j). However, low-dimensional dynamics also arise in the case (Figure 3k,l)
$V$ $=\left[\begin{array}[]{cc}-{\bf{u}}_{2}&{\bf{v}}_{\perp}\end{array}\right].$
in which case $W$ lacks eigenvalues with large, negative real part (Figure 3m). Hence, these two examples demonstrate that low-dimensional dynamics can arise in the presence or absence of eigenvalues with large, negative real part.
So far, in all of the examples with high-dimensional dynamics, $W_{0}$ is an EP matrix. However, $W_{0}$ need not be fully EP for dynamics to be high-dimensional. We now consider the case in which left and right singular vectors are not perfectly aligned, but have some non-vanishing overlap,
$V$ $=\sqrt{1-c}\left[\begin{array}[]{cc}-{\bf{u}}_{1}&\,{\bf{v}}_{\perp}\end{array}\right]+\sqrt{c}\left[\begin{array}[]{cc}{\bf{-}}{\bf{u}}_{1}&\,-{\bf{u}}_{2}\end{array}\right]$ (10)
where $0<c<1$. This is a linear combination of the EP and non-EP matrices from the examples above. In this case, $W_{0}$ is not EP, but $P$ has a singular value at $\sigma_{P}=\sqrt{c}$ and the other at $\sigma_{P}=1$ (Figure 3n). As long as $c$ is not close to zero, the condition in Claim 2 is satisfied, and the network produces high-dimensional dynamics (Figure 3o). Also, $W$ has eigenvalues with large, negative real part (Figure 3p). From this example, we may conclude that the low-rank part of the connectivity matrix only needs to have a non-vanishing EP component to produce high-dimensional dynamics.
The numerical results in Figure 3 provide an interpretation of Claim 2: If $W_{0}$ has a non-vanishing component with the EP property, then dynamics are high-dimensional. To our knowledge, this connection between the EP property of matrices and the dimensionality of network dynamics has not been made before.
The examples in Figure 3 also demonstrate that the eigenvalue spectrum is generally not useful for determining whether dynamics are high- or low-dimensional in strongly low-rank networks with high-dimensional input. As indicated above by Claim 1, the eigenvalue spectrum is also not useful for determining whether strongly low-rank networks produce low-rank suppression.
It is common in theoretical work to consider models in which the entries of $U$ and $V$ are random and independent with zero mean. In this case, ${\bf{u}}_{j}$ and ${\bf{v}}_{k}$ are nearly orthogonal so $\sigma_{P}\ll 1$ and the network produces low-dimensional dynamics (Figure 12 in Appendix D; compare to Figure 2). However, many networks arising in nature do not have purely random structure. In the following sections, we show that many naturally arising network structures satisfy our conditions for high-dimensional dynamics.
@@
If $m$ and $s$ scale similarly, then the largest singular value of $W$ is near $\sigma_{1}=|m|N$ while the next-largest singular value scales like $\sigma_{2}=\mathcal{O}(s\sqrt{N})$, implying an effective rank-one structure when $N$ is large. The dominant rank-one part has constant entries, so it is normal and the network exhibits low-rank suppression and high-dimensional dynamics (Figure 15 in Appendix D; see Appendix B in [35] for related results).
@@
where ${\bf{1}}_{e}=\{1,2,\ldots,N_{e}\}$ and ${\bf{1}}_{i}=\{N_{e}+1,N_{e}+2,\ldots,N\}$. Therefore, perturbations that are uniform within each population are aligned to the low-rank part and suppressed relative to random perturbations (Figure 4d,e). In other words, perturbations that stimulate all excitatory neurons equally and all inhibitory neurons equally are suppressed relative to perturbations that are inhomogeneous within one or both populations, consistent with previous work on balanced network models of cortical circuit responses [43, 23, 22, 24]. As predicted, simulations also show high-dimensional responses to high-dimensional inputs (Figure 16 in Appendix D), consistent with observations that neural responses in monkey visual cortex are high-dimensional when visual stimuli are high-dimensional [16].
@@
Many networks in nature exhibit connection strength that depends smoothly on the distance between nodes in physical or other spaces, resulting in an effective low-rank structure [1, 41]. As a simple example, we consider a model in which each node is assigned a spatial location, $\theta\in[0,1)$, and connection strength decays like a Guassian function of the distance, $d\theta$, between nodes so that only nearby nodes are strongly connected (Figure 5a). Connectivity is also perturbed by a random component, $W_{1}$, as above. This connectivity structure is effectively low-rank (Figure 5b) and the low-rank part is normal, with left- and right singular vectors spanned by the low-frequency Fourier basis vectors (Figure 17 in Appendix D). Therefore, networks of this form satisfy our conditions for low-rank suppression and high-dimensional dynamics (Figure 18 in Appendix D).
@@
where $g>0$, $\gamma>0$, and $\circ$ denotes element-wise multiplication. Each ${\bf{z}}_{j}(t)$ models the probability that an individual is infected. Here, $W$ is the proximity matrix of 637 high school students, indicating whether each pair of students were in proximity of each other during a specific school week [51]. This matrix is effectively low-rank in the sense that it has a small number of dominant singular values (Figure 19a in Appendix D). In [1], it was shown that the dynamics generated by Eq. (11) on this network are effectively low-dimensional.
@@
Simulations of these dynamics indeed demonstrated low-rank suppression (Figure 6a and Figure 19b in Appendix D) and high-dimensional dynamics (Figure 6b) in contrast to the model from [1]. These results highlight the importance of accounting for external perturbations and internal noise when studying the dimensionality of epidemiological dynamics. Moreover, the results imply that epidemiological networks are, perhaps counterintuitively, more sensitive to random perturbations than to perturbations aligned with the network’s low-rank structure.